Grade 12Algebra

Give first the step you will use to separate the variable and then solve the equation:
y – 4 = 4

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

We have to add 4 to both the side of given equation,
Then we get,
= y - 4 +4 = 4 + 4
=y =8

Thanks